# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Problem-Space Evasion Adversarial Attacks on Flow-based ML-NIDS

### v3

@@ 5.4 Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ5)
As shown in Table 5, the majority of perturbed or affected features in these problem-space evasion attacks exhibit high perturb-ability scores, classified under the red category. A smaller subset of features shows medium perturb-ability. This highlights the efficacy of our defense mechanism, which strategically eliminates high-perturb-ability features, rendering these types of attacks ineffective. Since the attacker manipulates features that are excluded from the final feature set. In other words, our defense will eliminate these easily perturbed features, which may render these types of attacks ineffective, as the attacker is modifying features that are excluded from the selection, making their changes irrelevant to the feature vector. Table 5 highlights a crucial observation: minimal modifications in the problem-space can lead to significant alterations across numerous features in the feature-space. While this may suggest that attackers possess considerable power or strategic advantage, these manipulations often introduce unintended side effects [23], which we term collateral damage. We coined the term collateral damage features because the term side effect features, commonly used in research, is primarily associated with the unintended consequences of medicines, which are generally beneficial. In contrast, within the context of adversarial attacks, these features are unintentionally perturbed by attackers, and adversarial attacks are generally detrimental. Therefore, we believe that collateral damage features is a more accurate and fitting term. In practical terms, when an attacker aims to modify a specific feature based on the gradient of the targeted mode, for example, the maximum forward IAT, they might induce a substantial delay between two packets in the flow. However, due to the inherent correlations between features in an ML-based NIDS, this single modification can unintentionally impact a broader set of features. For instance, changing the maximum forward IAT could also affect metrics such as flow duration, packets per second, total forward IAT, and forward IAT mean. This collateral damage does not follow any particular direction of the gradient [35], introducing an element of unpredictability. As a result, these unintentional feature perturbations may negatively influence the success of the attack, potentially undermining the adversary’s objectives. It is crucial to highlight that while most features can be modified in the problem-space, labeling a feature as non-perturb-able means it is difficult to manipulate while complying with NIDS domain constraints.
6 Related Work
@@
Beyond these general strategies, researchers have explored defenses that focus on feature manipulation for security applications. For instance, Chen et al. [14] introduced cost-aware training, incorporating domain-specific costs into defense strategies, where manipulating features incurs varying levels of difficulty (e.g., asymmetric costs for increasing or decreasing a feature). Unlike their approach, our method leverages the Perturb-ability Score (PS) as a generalizable metric for evaluating the robustness of NIDS features. PS enables the identification and selection of features that are inherently difficult to perturb, thereby reducing the attack surface without adding overhead to the training or testing phases. Furthermore, our mechanism is completely independent of the ML model and can be applied to any model, unlike the method in [14], which focuses specifically on tree ensemble models.
7 Discussion and Lessons Learned
7.1 The Usual Suspects
@@ 7.2 Five Features
Sheatsley et al. [34] argued that limiting the number of features an adversary can manipulate does not completely prevent adversarial attacks. Their findings demonstrated that adversarial samples could achieve a 50% success rate by modifying just five randomly selected features. We would like to point out that our PS-enabled defense does not aim to reduce the number of features but rather to select features that are non-perturb-able in the problem-space during the feature selection phase. Thus, the fundamental purpose of our suggested defense is to leave attackers with as few perturb-able features as possible (as seen in Table 5, the attackers will likely have access to far fewer than five perturb-able features). Moreover, The results by Sheatsley et al. [34] focused on targeting NIDS at the feature layer rather than the problem-space. Real-world network environments are far more complex, and the constraints modeled in their research, such as those related to protocols like TCP and UDP, are a small subset compared to the numerous limitations encountered in the actual problem-space, including side-effect features.
@@
7.3 Problem-space Evasion Adversarial Attacks are Already Extremely Hard for an Attacker
@@
Evasion adversarial attacks against ML-NIDS are highly complex and impractical for attackers [35]. ElShehaby et al. [35] identified key challenges, including limited attacker access to feature vectors, correlations between NIDS features, restricted knowledge of models and feature extraction, and the dynamic nature of modern ML-NIDS. Problem-space evasion attacks face additional hurdles, such as the Inverse Feature-Mapping Problem, which complicates the translation of feature-space perturbations into realistic network packet modifications [31], often causing problem-space attacks to fail to align with intended adversarial features. Moreover, some problem-space perturbations can sabotage attack efficacy, for example, adding delays as perturbations might reduce the effectiveness of DoS attacks.
@@
Thus, introducing a simple addition in the feature-selecting phase in the architecture of ML-NIDS through the usage of the PS scoring mechanism to eliminate easily perturb-able features could be the last nail in the coffin for these already highly impractical and complex problem-space evasion adversarial attacks against NIDS. Furthermore, in real-world adversarial attacks, guessing often plays a pivotal role [6]. In the NIDS domain, adversaries typically lack access to the attacked model’s internal information and cannot directly query the target system. As a result, they may need to predict the selected features of the target system [35]. If these selected features are unpredictable and difficult to perturb within the problem-space, attackers will be significantly constrained in crafting effective adversarial examples.
@@
7.4 Feature Reduction as a Defense Against Adversarial Attacks
@@
Feature squeezing and reduction have also been explored as defenses, narrowing the attacker’s search space [37, 25]. However, our approach goes beyond simple feature reduction by applying domain-specific constraints to select resilient features using PS. These features resist perturbations that would either compromise the network’s functionality or render the attack ineffective. By focusing on such features, our defense significantly reduces the effectiveness of problem-space adversarial attacks, making it much harder for attackers to succeed without disrupting network operations or their malicious intent.
7.5 NIDS Datasets
We recognize that existing NIDS datasets suffer from several limitations, including insufficient data diversity and ambiguities in ground truth labeling [18, 28]. To address these shortcomings, we employed the enhanced version of the CSE-CIC-IDS2018 dataset developed by Liu et al. [28]. Their work identified critical challenges in the original dataset, particularly in feature generation and labeling processes, which they rectified by modifying the CICFlowMeter tool and incorporating additional features.
7.6 We are losing information!
@@
Some may argue that utilizing PS to drop certain features could result in the loss of information that may be important for the ML-NIDS. While this concern is valid, it is essential that the application of PS be guided by architects, practitioners, and engineers with substantial domain knowledge. On the other hand, our results might suggest that current literature might be utilizing more information and features than necessary. Table 4 presents promising results with only the green features. Another pertinent question is: why rely on features that are easily perturb-able by an outsider?
7.7 Adversarial Attacks on ML-NIDS Research Direction
During our exploration for this work, we observed common issues in research on adversarial attacks against NIDS. Many studies are overly generous to the attacker, often assuming access to information that is rarely available in real-world settings. Grosse et al.[20] noted that this issue is common across artificial intelligence security applications research generally. In the NIDS domain, many research studies on evasion adversarial attacks presume full knowledge of the model and its selected feature set, which is not feasible in practice. Moreover, research on black-box evasion adversarial attacks assumes the possibility of querying ML-NIDS, which is likely to be infeasible in real-world scenarios [35]. As established in this paper, problem-space adversarial attacks are more practical than feature-space attacks. However, even for these attacks, existing evaluations tend to focus on their evasiveness rather than their maliciousness [23]. Additionally, research on problem-space adversarial attacks against ML-NIDS has largely overlooked the impact of side effects/collateral damage features [23]. There are also considerable issues within NIDS datasets that must be addressed [18].
8 Conclusion
This paper investigates the perturb-ability of features within NIDS by introducing the Perturb-ability Score (PS) metric. A high PS indicates that a feature is easily perturb-able in the problem-space while maintaining NIDS domain constrains, while a low PS suggests that altering that feature may either be infeasible or could render the network flow or malicious objective invalid. Through our research, we have leveraged the PS as a potential defense mechanism against practical problem-space evasion adversarial attacks by selectively identifying and utilizing features with low PS. This novel strategy introduces a fresh approach, utilizing NIDS domain problem-space constraints to mitigate the impact of adversarial attacks on Flow-based ML-NIDS. The results from our experiments demonstrate that discarding features with high PS does not compromise model performance. Instead, it enhances the robustness of NIDS against adversarial threats, providing reliable detection capability. The findings suggest that by focusing on inherently robust features, we can effectively fortify NIDS against problem-space evasion adversarial attacks while maintaining high detection accuracy. Overall, this work presents a promising avenue for enhancing the security of ML-based NIDS. Future research will focus on further refining the PS metric and exploring its applicability across different datasets and attack scenarios, ultimately contributing to the development of more resilient network security systems.
Appendix A Definition of Features of UNSW-NB15 and improved CSE-CIC-IDS2018 Datasets
Table 6: The Features Description of the UNSW-NB15 Dataset

### v4

@@ IX Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ4)
As shown in Table VIII, the majority of perturbed or affected features in these problem-space evasion attacks exhibit high perturb-ability scores, classified under the red category. A smaller subset of features shows medium perturb-ability. This highlights the efficacy of our defense mechanisms, which strategically eliminate or mask high-perturb-ability features. These defenses significantly reduce the effectiveness of such attacks, as the attacker manipulates features that are either masked or excluded entirely from the final feature set, making their modifications irrelevant to the feature vector used for classification.
Table VIII shows that limited changes in the problem-space can cause significant, widespread changes in the feature-space. While this may seem to benefit attackers, such manipulations often result in unintended side effects [25], which we term collateral damage. We coined the term “collateral damage features” because “side effect,” commonly used in research, is primarily associated with the unintended consequences of medicines, which are generally beneficial. In contrast, within the context of adversarial attacks, these features are unintentionally perturbed by attackers, and adversarial attacks are generally harmful. Therefore, we believe that ”collateral damage features” is a more accurate and fitting term. For example, modifying one feature (such as maximum forward IAT) to evade detection may unintentionally alter other correlated features (such as flow duration, packets per second, total forward IAT, and forward IAT mean) due to feature interdependencies in ML-based NIDS. These collateral changes do not follow any particular gradient direction [1], making their effects unpredictable and potentially undermining the attack.
X Adversarial Testing (RQ5)
X-A Systematization of ML-NIDS Adversarial Perturbations
Figure 11 presents a hierarchical classification of perturbation attack strategies against ML-NIDS, organized in four levels of increasing practicality and sophistication, with the innermost circle (Level 4) being the most practical and complex to create. The diagram uses overlapping circles to illustrate how each level builds upon and encompasses the constraints of previous levels:
Level 1 (Least Practical Attacks): Adding perturbations on features without taking NIDS features constraints into consideration. Level 2: Adding perturbations on features while taking NIDS features constraints (correlations, categorical and binary features) into consideration. Level 3: Adding perturbations on network traffic that produces perturbed features while taking into consideration NIDS features constraints and network functionality, but without maintaining the malicious aim of the attack after altering the network traffic. For example, adding huge delays between packets as perturbations in a A denial of service (DoS) attack. Level 4 (Most Practical Attacks): Adding perturbations on network traffic that produces perturbed features while taking into consideration NIDS features constraints, network functionality and malicious functionality.
In our tests to evaluate the effectiveness of our PS-enabled defenses, we utilized Levels 2 and 4, one in the feature-space and the other in the problem-space, to assess the robustness and diversity of our defenses. Moreover, level 2 provides a balance between practicality and coverage. The nested circular design of our illustration demonstrates how defending at a certain level will implicitly defend against higher levels.
Figure 11: ML-NIDS Adversarial Perturbations Types
Attack Success Rate (ASR): We evaluated the impact of adversarial attacks using the Attack Success Rate (ASR), defined as:
ASR=Number of Successful Adversarial AttemptsTotal Number of Adversarial $\text{ASR}=\frac{\text{Number of Successful Adversarial Attempts}}{\text{TotalNumber of Adversarial Attempts}}\times 100\%$
Prior research frequently defines a successful adversarial attack as occurring when the model’s prediction for a perturbed sample diverges from its original output:
$f(x_{i}^{*})\neq y_{i}$
Where $f(x_{i}^{*})$ represents the prediction for the perturbed sample $x_{i}^{*}$ and $y_{i}$ denotes the correct prediction. However, we argue this definition combines two distinct scenarios of prediction errors: (1) errors caused by adversarial perturbations and (2) inherent model inaccuracies on unperturbed inputs (the case when even the unperturbed sample produces wrong predictions).
To isolate the adversarial effect, we define a successful attack only when perturbations alter the model’s original correct prediction ($f(x_{i})=y_{i}$). Our revised ASR formulation reflects this distinction:
$\text{ASR}=\frac{\sum_{i=1}^{N}1((f(x_{i})=y_{i})\land(f(x_{i})\neq f(x_{i}^{*})))}{N}\times 100$
where $f(x_{i})$ denotes the model’s prediction for original input $x_{i}$, $f(x_{i}^{*})$ represents the perturbed input prediction, $N$ is the total number of adversarial attempts, and $1(\cdot)$ indicates the indicator function (1 if condition holds, 0 otherwise). The condition ($((f(x_{i})=y_{i})\land(f(x_{i})\neq f(x_{i}^{*})))$) ensures we only count instances where the original prediction was correct and the adversarial example successfully changed the prediction.
X-B Problem-space Attack vs Our Option A, B1 & B2 PS-enabled Defenses (level 4)
We conducted a series of experiments using the problem-space blind adversarial perturbation attacks proposed by Apruzzese et al. [18]. In our opinion, this is one of the most practical adversarial attacks in research. The authors assume that the attacker has no knowledge of the model, cannot query the model, and that their attack operates in problem-space (perturbing network flows) while maintaining NIDS constraints. Their blind perturbations manipulate UDP and TCP packets by adding a small padding of random bytes to their payloads. While applying these manipulations, they ensure that each packet does not exceed its maximum length and recreate the checksum. In our experiment, we used their published attack codes, the same PCAP dataset (MCFP dataset [30]) and the same tool to extract NetFlows from our raw PCAP traces (Argus [35]), with the same configuration as Apruzzese et al. [18]. We created four ML models for classification tasks and performed preprocessing to prepare the data for training.
To evaluate the impact of PS-guided feature selection and masking on the problem-space MCFP dataset, we extended our experiments from Section VIII. Table IX presents the performance metrics (Accuracy, Precision, Recall, F1-score) for four models (ANN, SVM, RF, CNN) under different defenses: pre-PS (baseline), post-PS feature selection (Option A), post-PS feature masking during training and inference (Option B1), and post-PS feature masking during inference only (Option B2). For both B1 and B2, we used the median value per feature from the training sets as the masking value for these experiments. Both Option A and Option B1 defenses preserved model performance. However, Option B2 resulted in noticeable performance drops in some models, particularly for SVM (cyan cells), and slight reductions in precision for RF and CNN (light blue cells). We will further discuss the limitations and ways to address them for Option B2 in Section XIII. For all experiments in this section, IP addresses were excluded during pre-processing to align with Apruzzese et al. [18], ensuring consistency.
To test the effectiveness of our PS-enabled defenses against the problem-space blind adversarial perturbation, the features were classified using PS to identify and exclude/mask easily perturbable features. The models were evaluated before and after applying PS-guided defenses. The results, shown in Table X, indicate that adversarial attacks by Apruzzese et al., while inconsistent, can reduce ML-NIDS performance. However, after applying PS-guided defenses (our Option A, B1, and B2), these attacks have no measurable effect on model performance. The results in Tables IX and X confirm that PS-guided defenses effectively mitigate adversarial attacks without compromising performance, except for inference-only masking in Option B2.
TABLE IX: Performance metrics (Accuracy, Precision, Recall, F1-score) for models before and after PS feature selection on MCFP Dataset
Model Acc Pre Rec F1
Pre-PS (Baseline) ANN 0.9998 1.0000 0.9990 0.9995
SVM 0.9950 0.9736 0.9970 0.9852
RF 1.0000 1.0000 1.0000 1.0000
CNN 0.9997 0.9980 1.0000 0.9990
  Post-PS feature selection (Only (Green) features Selected) Option A Defense ANN 0.9997 0.9990 0.9990 0.9990
SVM 0.9963 0.9804 0.9980 0.9891
RF 1.0000 1.0000 1.0000 1.0000
CNN 1.0000 1.0000 1.0000 1.0000
  Post-PS feature Masking in Training and Inference (Only (Green)
features are not masked)
Option B1 Defense
ANN 0.9998 1.0000 0.9990 0.9995
SVM 0.9947 0.9717 0.9970 0.9842
RF 1.0000 1.0000 1.0000 1.0000
CNN 1.0000 1.0000 1.0000 1.0000
  Post-PS feature Masking in Inference (Only (Green) features are not masked) Option B2 Defense ANN 0.9932 0.9651 0.9950 0.9798
SVM 0.8710 0.7927 0.3060 0.4416
RF 0.9842 0.9132 1.0000 0.9547
CNN 0.9792 0.8889 1.0000 0.9412
TABLE X: ASR Results for TCP and UDP with Pre PS-enabled and Post PS-enabled Defenses (Options A, B1 & B2) on MCFP Dataset
Model TCP ASR UDP ASR
Pre PS-enabled Defenses ANN 0.00% 0.00%
SVM 22.50% 39.40%
RF 0.00% 0.00%
CNN 0.00% 38.40%
  Post Option A Defense: Only (Green) features Selected Options B1 & B2 showed the exact results of 0% ASR across all models. In Options B1&2: only (Green) features weren’t masked. ANN 0.00% 0.00%
SVM 0.00% 0.00%
RF 0.00% 0.00%
CNN 0.00% 0.00%
X-C Constrained Feature-space Attack vs Our Option B2 PS-enabled Defense (level 2)
To test the effectiveness of our option B2 defense, we conducted a series of experiments using the work by Alhussien et al. [36]. They introduced a comprehensive set of network constraints specifically designed to ensure the validity and practicality of adversarial examples in network traffic scenarios. They evaluated their approach using multiple attacks like; Zeroth Order Optimization (ZOO) [37], a black-box attack relying on gradient estimation through model queries, and Carlini & Wagner (C&W) [19], a white-box optimization-based attack known for producing minimal perturbations. Additionally, Alhussien et al. published portions of their source code, allowing us to utilize and extend their constrained adversarial attack generation methodologies within our experimental evaluation. To reproduce Alhussien et al.’s results and tests, we used the same models, dataset, pre-processing, feature selection, attacks, attack settings, and constraints provided in their published codes. To ensure consistency and avoid altering their architecture or pipeline, we employed Option B2 (inference-only masking), which allowed us to dynamically mask high perturb-ability features during inference while maintaining the original model structure and pre-trained parameters. For more details on Alhussien et al.’s [36] constrained attacks and our testing of our defenses against these attacks, including the constraints they imposed on the attacks, the models and attacks we used, and limitations, check Appendix B.
TABLE XI: Comparison of Constrained ZOO and C&W Attacks Pre PS-enabled and Post PS-enabled Feature Masking (Our Option B2 Defense) on UNSW-NB15 Dataset
Attack Constrained ZOO Constrained C&W
Adversarial Test Samples 4533 4533
Successful Samples 897 2706
ASR Pre PS-enabled Masking 19.79% 59.70%
  ASR Post PS-enabled Masking red + yellow (Only Keeping green not masked) 0.00% 0.00%
ASR Post PS-enabled Masking red (Keeping green and yellow not masked) 0.00% 0.00%
Table XI summarizes our experimental results, comparing the effectiveness of these constrained attacks before and after applying our PS-enabled inference feature masking defense (Option B2). The masking in these tests was applied by masking with the neutral value of 0.5 as Alhussien et al. used min-max normalization. Initially, the constrained ZOO attack achieved an Attack Success Rate (ASR) of 19.79%, while the constrained C&W attack attained an ASR of 59.70%. However, after applying our PS-enabled feature masking, both when masking red and yellow features (keeping only green features unmasked) and when masking only red features (keeping green and yellow unmasked), the ASR dropped to 0.00% for both attacks. These results demonstrate the effectiveness of our proposed PS-enabled defense in mitigating constrained adversarial attacks on flow-based ML-NIDS. Notably, the performance of both models remained consistent in these sets of experiments, maintaining near-perfect F1 and accuracy scores ($>$0.99) after applying our B2 defense.
X-D Adversarial Training
To further evaluate the effectiveness of our PS-enabled defenses, we performed adversarial training experiments. Table XII summarizes the results of adversarial training on the UNSW-NB15 dataset using constrained ZOO and C&W attacks. We generated 12,666 adversarial training samples from a total of 27,544 training samples. The constrained ZOO attack required approximately 40 minutes, while the constrained C&W attack took around 44 minutes to generate these samples. After adversarial training, the Attack Success Rate (ASR) significantly decreased to 0.81% for constrained ZOO and 3.89% for constrained C&W attacks.
Although adversarial training is widely regarded as one of the most effective approaches to defending against adversarial examples [38], and as seen in Table XII, it reduces the ASR of the attacks, it has several drawbacks compared to our PS-enabled defenses. First, adversarial training is not universal; its effectiveness depends on incorporating specific attack types into the training process, making the model robust mainly against these types of attacks. As highlighted by Abou Khamis et al. [39], finding a general adversarial training defense for ML-based IDS remains challenging. In contrast, our PS-enabled defense is universal (attack independent), leveraging problem-space constraints inherent to NIDSs to effectively ”tie the hands” of attackers without requiring retraining or attack-specific adjustments. Furthermore, our Option B defense is extremely lightweight (it has no time overhead) compared to adversarial training, which demands significant computational resources for generating adversarial samples and retraining models (as evidenced by the time metrics in Table XII. In summary, while adversarial training provides robust protection against specific attacks, its inefficiency and lack of universality make it less practical than our PS-enabled defense for real-world applications.
TABLE XII: Effect of Adversarial Training on Constrained ZOO and C&W Attacks on UNSW-NB15 Dataset
Attack Constrained ZOO Constrained C&W
Total Training Samples 27544 27544
Adversarial Training Samples 12666 12666
Time to generate The Adversarial Training Samples (in minutes:seconds) 40:23 44:10
Adversarial Test Samples 2110 2110
Successful Samples 17 82
ASR After Adversarial Training 0.81% 3.89%
XI Related Work
@@
Beyond these general strategies, researchers have explored defenses that focus on feature manipulation for security applications. For instance, Chen et al. [44] introduced cost-aware training, incorporating domain-specific costs into defense strategies, where manipulating features incurs varying levels of difficulty (e.g., asymmetric costs for increasing or decreasing a feature). Unlike their approach, our method leverages the PS metric as a generalizable measure for evaluating the perturb-ability of NIDS features in problem-space. PS enables the identification and elimination or masking of features that are inherently easy to perturb, thereby reducing the attack surface without adding overhead to the training or inference phases. Furthermore, our mechanism is completely independent of the ML model and can be applied to any model, unlike the method in [44], which focuses specifically on tree ensemble models.
XII Discussion and Lessons Learned
XII-A The Usual Suspects
@@ XII-B Five Features
Sheatsley et al. [24] argued that limiting the number of features an adversary can manipulate does not completely prevent adversarial attacks. Their findings demonstrated that adversarial samples could achieve a 50% success rate by modifying just five randomly selected features. We would like to point out that our PS-enabled defenses do not aim to reduce the number of features but rather to select non-perturb-able features during the feature selection phase or mask perturb-able features in the problem-space. Thus, the fundamental purpose of our suggested defenses is to leave attackers with as few perturb-able features as possible (as seen in Table VIII, the attackers will probably have access to fewer than five perturb-able features). Moreover, the results by Sheatsley et al. [24] focused on targeting NIDS at the feature layer rather than the problem-space. Real-world network environments are far more complex, and the constraints modeled in their research, such as those related to protocols like TCP and UDP, are a small subset compared to the numerous limitations encountered in the actual problem-space, including side-effect features.
@@
XII-C Problem-space Evasion Adversarial Attacks are Already Extremely Hard for an Attacker
@@
Evasion adversarial attacks against ML-NIDS are highly complex and impractical for attackers [1]. ElShehaby et al. [1] identified key challenges, including limited attacker access to feature vectors, correlations between NIDS features, restricted knowledge of models and feature extraction, and the dynamic nature of modern ML-NIDS. Problem-space evasion attacks face additional hurdles, such as the Inverse Feature-Mapping Problem, which complicates the translation of feature-space perturbations into realistic network packet modifications [8], often causing problem-space attacks to fail to align with intended adversarial features. Moreover, some problem-space perturbations can sabotage attack efficacy; for example, adding delays as perturbations might reduce the effectiveness of DoS attacks. However, there are rare glimpses of practical evasion adversarial attacks against flow-based NIDS in research, such as the “blind” adversarial perturbations in [18].
@@
Thus, introducing a simple addition in the feature-selecting phase of ML-NIDS architecture through the PS scoring mechanism to eliminate or mask easily perturb-able features could decisively undermine these already highly impractical and complex problem-space evasion attacks. Furthermore, in real-world adversarial attacks, guessing often plays a pivotal role [10]. In the NIDS domain, adversaries typically lack access to the attacked model’s internal information and cannot directly query the target system. As a result, they may need to predict the selected features of the target system [1]. If these selected features are unpredictable and difficult to perturb within the problem-space, attackers will be significantly constrained in crafting effective adversarial examples.
@@
XII-D Feature Reduction and Masking as Defenses Against Evasion Adversarial Attacks
@@
Feature squeezing and reduction have been explored as defenses, narrowing the attacker’s search space [13, 3]. However, our approach goes beyond simple feature reduction by applying domain-specific constraints to select resilient features using PS during feature selection (our Option A Defense). These features resist perturbations that would either compromise the network’s functionality or render the attack ineffective. Furthermore, adding random padding [45] or noise [46] or masking specific pixels [47] has been explored as defenses against evasion adversarial attacks. However, to the best of our knowledge, our PS-enabled feature masking is the first to mask specific features based on their susceptibility to problem-space perturbations in the flow-based ML-NIDS domain. By focusing on resilient features, our defense strategies significantly reduce the effectiveness of problem-space adversarial attacks on flow-based ML-NIDS, making it much harder for attackers to succeed without disrupting network operations or their malicious objectives.
XII-E We are losing information!
@@
Some may argue that utilizing PS to drop or mask certain features could result in the loss of information that may be important for the ML-NIDS. While this concern is valid, it is essential that the application of PS be guided by architects, practitioners, and engineers with substantial domain knowledge. On the other hand, our results might suggest that current literature might be utilizing more information and features than necessary. Tables V, VI and VII present promising results when only low-perturb features are selected or when high-perturb features are masked. Another pertinent question is: why rely on features that are easily perturb-able by an outsider?
XIII Limitations
Although our PS-enabled defenses are equally effective in eliminating the impact of evasion adversarial attacks against ML-NIDS, they are evasion-attack-agnostic, model-agnostic, and lighter than legacy approaches such as adversarial training; however, their stability varies across implementations. Options A and B1 demonstrate robust and stable performance by integrating PS-guided feature selection or masking during training. Option B2, while the lightest defense (requiring no retraining or architectural changes) and equally effective at neutralizing attacks, exhibits instability in some scenarios, as seen in Table IX (e.g., SVM’s F1-score drop to 0.44 in cyan cells). We acknowledge that because masking in Option B2 is applied only during inference, it may mask features that the model relies on for accurate predictions, leading to performance degradation in certain cases.
However, post-processing techniques might mitigate these performance issues. For instance, applying class reweighting, different/dynamic masking values, probability calibration, and F1-optimized thresholding might significantly improve results for vulnerable models like SVM. We applied a combination of these mitigations, and this increased the SVM’s F1-score from 0.44 (in Table IX) to 0.96 for class 0 (benign traffic) and 0.82 for class 1 (malicious traffic), with a weighted average F1 of 0.93. By calibrating probabilities to reflect true class distributions and dynamically selecting thresholds that balance precision and recall, we restore performance without compromising the defense’s lightweight nature. Thus, while Option B2 may require supplemental post-processing for unstable edge cases (a tradeoff we plan to investigate further) it remains a viable on-the-fly defense for resource-constrained deployments in some cases.
XIV Conclusion
The PS metric provides a novel approach to enhancing the robustness of flow-based Machine Learning (ML)-NIDS against problem-space evasion adversarial attacks. We proposed two PS-enabled defense mechanisms: PS-guided feature selection, which utilizes only features with low PS during the feature selection phase, and PS-guided feature masking, which masks high PS features with neutral values. Both approaches leverage NIDS domain problem-space constraints to mitigate the impact of adversarial attacks on flow-based ML-NIDS. Our experimental results demonstrate that these PS-enabled defenses significantly enhance the robustness of ML-NIDS against adversarial threats while maintaining high detection accuracy.
The PS metric and associated defenses offer several key advantages: universal applicability across various evasion adversarial attacks, ML models and datasets in the flow-based NIDS domain, lightweight implementation, practicality in addressing real-world attack scenarios, and complementarity to existing defenses. This work opens up promising avenues for future research, including refinement of the PS metric, cross-domain application, dynamic PS adaptation, and integration with advanced ML techniques. Overall, this research presents a significant advancement in the field of ML-NIDS security, providing a practical, efficient, lightweight, and effective approach to enhancing the robustness of flow-based ML-NIDS against problem-space evasion adversarial attacks.
Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the NSERC Discovery Grant program.
Appendix A Ethics Considerations
The datasets used in this study, MCFP, UNSW-NB15 and the improved CSE-CIC-IDS2018, are publicly available and widely used in network intrusion detection research. Additionally, the proposed Perturb-ability Score (PS) metric was developed and evaluated using synthetic experiments that do not involve human participants.
Appendix B Additional Information on the constrained attacks by Alhussien et al. [36]
B-A Constraints imposed by Alhussien et al. [36] on adversarial attacks against NIDS
These constraints are categorized into four types: feature mutability constraints, which identify modifiable and immutable network features; feature value constraints, which enforce valid ranges or categories for features; feature dependency constraints, which preserve semantic and statistical relationships between interdependent features; and distribution-preserving constraints, which maintain the statistical distribution of original network data. To incorporate these constraints into adversarial attack generation, they augmented the optimization objective with constraint-based penalty terms.
B-B Details of the attacks (Constrained ZOO and C&W Attacks) and the Attacked Models
To test our defenses against the constrained attacks by Alhussien et al. [36] we employed two constrained adversarial attacks: the Zeroth Order Optimization (ZOO) attack and the Carlini & Wagner (C&W) attack. Both attacks were configured with an attack confidence of 0.5, a maximum of 80 iterations, and were executed as untargeted attacks. The constrained ZOO attack targeted a Gradient gradient-boosting classifier consisting of an ensemble model with 100 estimators, a learning rate of 1.0, a maximum depth of 5, and a random state set to 0. This model was trained on a balanced dataset. The constrained C&W attack targeted a Neural Network model with a three-layer architecture ([256]-[256]-[128]) employing ReLU activation functions and an Adam optimizer (learning rate = 0.01), trained for 100 epochs with early stopping based on test accuracy. Table XIII summarizes the details of the attacks and attacked models. N.B., we obtained these parameters and models from the published code of Alhussien et al. [36].
B-C Limitations
While we highly appreciate the work done by Alhussien et al. [36], we observed some limitations in their testing and code implementations. However, in our opinion, the main limitation is their focus on the feature-space in their testing rather than the problem-space. We acknowledge that their constrained attack captures many NIDS constraints, but the complexity of real-world network environments significantly exceeds the constraints modeled in the feature layer. Even in their results, after adding perturbations from their constrained attacks, 25% of C&W perturbed samples failed to mimic valid network traffic. Nevertheless, we utilized their work to evaluate our PS-enabled defense. As illustrated in Figure 11, we believe that defending against constrained feature-space attacks, even if not entirely practical or valid in the problem-space, will implicitly defend against more practical problem-space attacks. In other words, even if an attack is not completely valid in the problem-space, our defense will still be effective, demonstrating the robustness of our PS-enabled approach.
TABLE XIII: Details of the attacks (Constrained ZOO and C&W Attacks) and the Attacked Models
Attack Constrained ZOO Constrained C&W
Attack Details Attack Confidence: 0.5
Max Iterations: 80
Untargeted
Attack Confidence: 0.5
Max Iterations: 80
Untargeted
Attacked ML Model Gradient Boosting Classifier Neural Network
Attacked ML Model Details Ensemble model with 100 estimators, learning rate=1.0, max_depth=5, random_state=0. Trained on balanced dataset. 3-layer architecture (input→[256][256][128] →2 outputs) with ReLU activation. Adam optimizer (lr=0.01), trained for 100 epochs with early stopping based on test accuracy.
Appendix C Definition of Features of UNSW-NB15 and improved CSE-CIC-IDS2018 Datasets
TABLE XIV: The Features Description of the UNSW-NB15 Dataset